3. Lindon Company is the exclusive distributor for an automotive product that sells for $18.00 per...

3.

Lindon Company is the exclusive distributor for an automotive product that sells for $18.00 per unit and has a CM ratio of 30%. The company’s fixed expenses are $92,340 per year. The company plans to sell 13,200 units this year.


Required:

1.

What are the variable expenses per unit? (Round your answer to 2 decimal places.)

     

2. Use the equation method:
a.

What is the break-even point in unit sales and in dollar sales?

            

b.

What amount of unit sales and dollar sales is required to earn an annual profit of $27,000?

          

c.

Assume that by using a more efficient shipper, the company is able to reduce its variable expenses by $1.80 per unit. What is the company’s new break-even point in unit sales and in dollar sales?

          

3. Repeat (2) above using the formula method.
a.

What is the break-even point in unit sales and in dollar sales?

            

b.

What amount of unit sales and dollar sales is required to earn an annual profit of $27,000?

          

c.

Assume that by using a more efficient shipper, the company is able to reduce its variable expenses by $1.80 per unit. What is the company’s new break-even point in unit sales and in dollar sales?
Answer #1
1) Variable expense per unit
18*70%
12.6
2a) (18.-12.6)=92340
17100 units
x *30% = 92340
x= 307800 dollar sales
2b) (5.4x) = 92340+27000
x= (92340+27000)/5.4
22100 units
dollar sales
x *30%= 92340+27000
x= 397800 answer
c) 7.2 x = 92340
x= 12825 units
x *40%= 92340
x = 230850
3a) BEP(units) = fixed cost/contribution per unit
92340/5.4
17100
BEP(dollar sales) = fixed cost/contribution margin
92340/30%
307800
b) unit sales = (fixed cost+target profit)/contribution per unit
(92340+27000)/5.4
22100 units
dollar sales = 22100*18
397800
c) new contribuiton = 5.4+1.8
7.2
BEP(units) = fixed cost/contribution per unit
92340/7.2
12825 answer
BEP(dollar sales) = 12825*18
230850 answer